Calculate the gross and net calorific values of coal having the following composition. C=83,H=7%,S=2%,N=4%,O=3%,Ash=1%
1
Expert's answer
2021-10-04T05:11:10-0400

N(C)=83/12=7

N(H)=7/1=7

N(S)=2/32=0.0625

N(O)=3/16=0.1875